OpenStudy (anonymous):

solve by using substitution? -5x+4y=20 and 10x-8y=-40 explain how to do these please, i have no idea :(

OpenStudy (shamil98):

-5x + 4y = 20 10x - 8y = -40 solve for y.. 4y = 5x + 20 y = 5/4 x + 5 now substitute into the other equation 10x - 8 (5/4x+5)= -40 and solve for x.